# SPA-Bench: A Comprehensive Benchmark for SmartPhone Agent Evaluation

### v1

@@ F.1 Agent Configuration
  • AppAgent, SeeAct, M3A, and T3A: Added AdbKeyboard [^6] https://github.com/senzhk/ADBKeyBoard for Chinese character input, following the MobileAgent setup.
@@ F.2 Experimental Results
Agent Success (%) Mean Step Ratio on Success Termination Reason Termination Inaccuracy Mean Exec Time per Step (sec) Mean Token Cost per Step (USD)
SRC (%) MSR (%) Error (%) Premature (%) Overdue (%)
Off-the-Shelf Model (GPT-4o)
@@
Agent Success (%) Mean Step Ratio on Success Termination Reason Termination Inaccuracy Mean Exec Time per Step (sec) Mean Token Cost per Step (USD)
SRC (%) MSR (%) Error (%) Premature (%) Overdue (%)
Off-the-Shelf Model (GPT-4o)
@@
Agent Success (%) Mean Step Ratio on Success Termination Reason Termination Inaccuracy Mean Exec Time per Step (sec) Mean Token Cost per Step (USD)
SRC (%) MSR (%) Error (%) Premature (%) Overdue (%)
Off-the-Shelf Model (GPT-4o)
@@ F.3 Performance Across Task Difficulty Levels
Table 14: Success rates on single-app English, single-app Chinese, cross-app English and cross-app Chinese tasks, categorised by difficulty level.
@@
F.4 Performance Loss Using Open-source Model
We explored how agents relying on proprietary (M)LLMs perform when replaced with advanced open-source models. Four agents (AutoDroid, AppAgent, MobileAgentV2, and M3A) were tested on Level 1 single-app tasks using open-source (M)LLMs based on their input modalities: Llama-3.1-8B-Instruct (no image), GLM-4V-9B (single image), and MiniCPM-V-2.6-8B (multiple images).
Table 15 compares performance with proprietary models (GPT-4o). M3A, which achieved a success rate of 0.800 with GPT-4o in level 1 single-app English tasks, dropped to 0.040 when using MiniCPM-V-2.6-8B. Similarly, MobileAgentV2’s success rate fell from 0.7 to 0. These results show that MiniCPM-V-2.6-8B performs poorly in smartphone control tasks, where long-sequence decision-making is critical. The error rate for MobileAgentV2 with MiniCPM-V-2.6-8B in English tasks reached 0.820 due to its occasional failure to follow instructions. In contrast, M3A’s error rate is only 0.060, as it converts invalid actions into a “wait” command, taking no action for that step, while MobileAgentV2 terminates upon encountering errors.
AutoDroid and AppAgent also showed similar trends, where proprietary models outperformed their open-source counterparts. However, since AutoDroid and AppAgent handle simpler input modalities, they performed relatively better and more stable with open-source models. Similar patterns were observed in Chinese tasks.
Table 15: Comparison of success and error rates across four agents based on GPT-4o and open-source models for Level 1 single-app English and Chinese tasks. Open-source model choices are determined by input modality.
Agent Model English Tasks Chinese Tasks
Success Rate Error Rate Success Rate Error Rate
AutoDroid GPT-4o 0.560 0.060 0.360 0.020
Llama-3.1-8B-Instruct 0.220 0.040 0.020 0
AppAgent GPT-4o 0.540 0.100 0.400 0.440
GLM-4V-9B 0.180 0.440 0.100 0.500
MobileAgentV2 GPT-4o 0.700 0 0.580 0.020
MiniCPM-V-2.6-8B 0 0.820 0.020 0.660
M3A GPT-4o 0.800 0 0.500 0
MiniCPM-V-2.6-8B 0.040 0.060 0.100 0
Appendix G Case Study
G.1 AppAgent on contact_2 task
@@
G.2 MobileAgent on expedia_3 task
@@
G.3 M3A vs Human on google_tasks_0 task
@@
By comparing Figure 15(a) and Figure 15(b), it is evident that M3A employed a shortcut strategy by combining text input and pressing the enter key into a single-step operation. This approach resulted in a more concise execution, with one fewer step compared to the human’s.

### v2

@@ F.1 Agent Configuration
  • AppAgent, SeeAct, M3A, and T3A: Added AdbKeyboardhttps://github.com/senzhk/ADBKeyBoard for Chinese character input, following the MobileAgent setup.
@@ F.2 Experimental Results
Agent Success Rate Mean Step Ratio on Success Termination Reason Termination Inaccuracy Mean Exec Time per Step (sec) Mean Token Cost per Step (USD)
SRC Rate MSR Rate Error Rate Premature Rate Overdue Rate
Agentic Workflow (GPT-4o)
@@
Agent Success Rate Mean Step Ratio on Success Termination Reason Termination Inaccuracy Mean Exec Time per Step (sec) Mean Token Cost per Step (USD)
SRC Rate MSR Rate Error Rate Premature Rate Overdue Rate
Agentic Workflow (GPT-4o)
@@
Agent Success Rate Mean Step Ratio on Success Termination Reason Termination Inaccuracy Mean Exec Time per Step (sec) Mean Token Cost per Step (USD)
SRC Rate MSR Rate Error Rate Premature Rate Overdue Rate
Agentic Workflow (GPT-4o)
@@ F.3 Performance Across Task Difficulty Levels
Table 14: Success rates on single-app English, single-app Chinese, cross-app English and cross-app Chinese tasks, categorised by difficulty level. AutoDroid was tested only on single-app tasks as its agent framework, Droidbot (Li et al., 2017), supports only these tasks.
@@
Appendix G Experiments on Open-ended Single-App English Tasks
To further explore the scalability of our success detection approaches, we designed an initial set of ten “open-ended” single-app English tasks across distinct apps, as detailed in Table LABEL:tab:english-single-tasks-2.
Table 15: Open-ended single-app English tasks.
App Task Description
Airbnb I’m traveling to London with three friends and need accommodation. I’$l$ manage the checkout process myself.
Amazon I’d like to buy wedding gifts for my friend and their partner. I’$l$ take care of the checkout myself.
Calculator I want to show my friend the multiplication of two negative numbers is indeed a positive number.
Chrome I’m planning a trip to ski and would like to save a blog to read later.
Clock Please set two alarms, one for weekdays and another for weekends. I prefer waking up later on weekends.
Merriam-Webster I’d like to expand my vocabulary in political ideologies. I aim to learn two new terms today.
Google Maps My car is low on gas, and I’m also feeling hungry.
Settings Sometimes I have trouble reading the screen clearly.
Spotify Create a music playlist for me in a recommended genre. Just two songs will do.
YouTube I’m interested in watching tech tutorial videos recently.
As discussed in Section 3.2, when a task description is clearly defined with a specific goal, its executions typically converge to the same final state. Such tasks can be treated as “closed-ended” tasks, which form the basis for human-annotated key components. In contrast, for a more vague task description, the task is considered “open-ended”. The final state may result in multiple possible outcomes, making it challenging to define key components explicitly. While the coarse detection phase may be limited in such cases, we hypothesised that our fine detection approach, relying on the MLLM evaluator, remains effective and can still be applied to “open-ended” tasks.
Table 16: Success rates on open-ended single-app English tasks.
Agent
Success
Rate
Agentic Workflow (GPT-4o)
AppAgent 0.200
AutoDroid 0.300
MobileAgent 0.200
MobileAgentV2 0.200
M3A 0.700
T3A 0.400
SeeAct 0.400
In this initial experiment, we tested the seven agents that follow the agentic workflow on the ten “open-ended” tasks. Given the open-ended nature of these tasks and the absence of predefined golden steps, agents were allowed a maximum of 20 steps to complete each task. We compared the alignment of success detection results between human evaluations and our MLLM evaluator. Using the same MLLM evaluator introduced in Section 5.2, we identified 22 true positives, 2 false positives, 2 false negatives, and 44 true negatives. This resulted in an F1 score of 0.917, consistent with the corresponding results for “closed-ended” tasks reported in Table 9. These findings demonstrate the potential of applying our MLLM evaluator to a broader range of tasks, both “open-ended” and “close-ended”, highlighting its scalability to tasks beyond our benchmark.
Table 16 presents the success rates of the seven agents on these ten tasks. M3A consistently outperformed the other agents. However, compared to the success rates reported in Table 3, MobileAgentV2 exhibited the largest performance gap, suggesting its limitations in handling “open-ended” tasks.
In future work, we aim to expand this initial experiment with a more comprehensive task collection to further improve and assess the feasibility of our MLLM evaluator for a wider range of tasks, and to investigate agent performance on “open-ended” tasks.
Appendix H Case Study
Three case studies are presented to illustrate representative scenarios of task execution by agents. These include: (1) an invalid action taken by AppAgent due to misinterpretation of the UI structure in the XML file, (2) a dynamically changing screen without any action execution, repetitive actions due to the lack of reflection, and unrelated behaviours to the task description in MobileAgent, and (3) the combined actions employed by M3A.
H.1 AppAgent on contact_2 task
@@
H.2 MobileAgent on expedia_3 task
@@
H.3 M3A vs Human on google_tasks_0 task
@@
By comparing Figure 15(a) and Figure 15(b), it is evident that M3A employed a combined action strategy, encapsulating text input and pressing the “enter” key within a single-step operation. This approach led to a more concise execution, requiring one fewer step compared to the human trajectory.